You have an aged relative who is terrified of debt. She runs a company that has zero debt. The company is in steady state. Its annual revenue is fixed at $1,000 and all income is distributed as dividends. There are 40 shares are outstanding. The dividends are $6.00 per share. Its required rate of return (or cost of equity) is 12%. With zero dividend growth this results in a current share price of $50.00. Currently the risk-free rate is 5% and the market risk premium is 7%. So the company’s stock beta is 1.00. This is also the asset beta.

Your aged relative retires and a new CEO is hired. The new CEO sells $1,000 of 7% 100-year bonds and repurchases 20 shares of stock. Nothing else changes in the company. The company’s tax rate is 40%. Construct new financial statements that reflect this change and estimate how much the market price of the stock will change, if at all. There will be no change in revenue and all income will be distributed as dividends.

Pre-Change Income Statement   

Post-Change Income Statement

Revenue

$1,000

$1,000

Expenses

600

Taxable Income

400

Taxes

160

Net Income

240

Dividends per share

$6.00

Share Price

$50.00

Pre change Balance Sheet

Post-Change Balance Sheet

Total Assets

$2,000

Debt

0

$1,000

Market Value of Equity

$2,000

Total Liabilities & Equity

$2,000
Answer #1

Income Statement   

Pre-Change Post-Change
Revenue 1000 1000
Expenses 600 670 (600+(1000*7%))
Taxable Income 400 330
Taxes 160 132
EPS=6 Net Income 240 198 EPS=198/20=9.9
240/40=6 Dividends per share 6 9.9 198/(40-20)
6/12%=50 Share Price 50 82.47 P/E=8.33
P/E=50/6=8.33 P/E=P/9.9=8.33
Balance Sheet Pre-Change Post-Change
Total Assets 2000 2649
Debt 0 1000
50*40 Market Value of Equity 2000 1649 20*82.47
Total Liabilities & Equity 2000 2649
So, share price increases to $ 82.47